# Computing Wasserstein Barycenters through Gradient Flows

### v2

@@ 3.4 Convergence
One of the difficulties in analyzing the gradient flow of the functional in equation 14 comes from the fact that $P\mapsto\mathbb{W}_{p}(P,Q)^{p}$ is not geodesically convex in $\mathcal{P}_{p,ac}(\Omega)$ [santambrogio2017euclidean, Section 4.4]. As a result, we are minimizing a non-convex functional. Henceforth, we focus on $p=2$, as this is the setting where most results are available. Explicit constants and proofs are available in Section 4 in the Appendix.
Our analysis relies on a measure-theoretic version of the Polyak Łojasiewicz (PL) inequality [polyak1964gradient, karimi2016linear],
$\lVert\nabla\mkern-10.0mu\nabla\mathbb{B}(P)\rVert_{L_{2}(P)}\geq C_{\text{PL}}(\mathbb{B}(P)-\mathbb{B}^{\star}),$ (22)
which is covered in [chewi2024statistical]. Under this assumption, we have the following convergence result,
@@
Let $\Omega=\mathcal{B}(0,R)$ be the closed ball on $\mathbb{R}^{d}$ with radius $R>0$. Let $P^{\star}\in\mathcal{P}_{2,\text{ac}}(\mathcal{B}(0,R))$ be the barycenter of $\mathcal{Q}=\{Q_{k}\}_{k=1}^{K},Q_{k}\in\mathcal{P}_{2,\text{ac}}(\mathcal{B}(0,R))$ $k=1,\cdots,K$, with barycentric coordinates $\lambda=(\lambda_{1},\cdots,\lambda_{K})\in\Delta_{K}$. Assume the inequality 47. Then, the following holds,
$\begin{aligned} \mathbb{E}[\hat{\mathbb{B}}(P_{\tau})-\hat{\mathbb{B}}^{\star}]\leq e^{-C_{\text{PL}}\tau}(\mathbb{B}(P_{0})-\mathbb{B}^{\star})+C_{R}\sqrt{\dfrac{C_{d,m}}{n}}\end{aligned},$ (23)
where the expectation on the l.h.s. is taken with respect to samples from $Q_{k}$. The constants $C_{R}$ and $C_{d,m}$ depend on the radius $R$, number of barycenter samples $m$, and dimensions $d$.
The right hand side of inequality 23 is composed of two terms. The first term comes from the PL inequality, and covers the convergence towards the minimizer of $\mathbb{B}$. The second term covers the error of empirical approximations. In this sense, letting $\tau\rightarrow+\infty$ leads to an error governed by the empirical approximation.
Next, we derive a new result for the convergence of the GMM gradient flow (cf. Algorithm 2). Here, we assume $P$ and $Q_{k}\in\mathcal{Q}$ are GMMs. Furthermore, we denote by $\hat{Q}_{k}$ the GMM fitted on data from $Q_{k}$ with the expectation-maximization algorithm. Our result relies on [boite2025differentiable, Theorem 3.1]. Therefore, we assume,
  1. 1.
    bounded GMM parameters, i.e., $\lVert\mu_{i}^{(Q_{k})}\lVert_{2}\leq R_{\mu}$, $\sqrt{\text{Tr}(\Sigma_{i}^{(Q_{k})})}\leq R_{\Sigma}$ $\forall i,\forall k$ (resp. $\hat{\mu}_{i}^{(Q_{k})}$ and $\hat{\Sigma}_{i}^{(Q_{k})}$, the estimated parameters); furthermore,
  2. 2.
    Empirical convergence rates, i.e., $\mathbb{E}[\lVert\pi^{(Q_{k})}-\hat{\pi}^{(Q_{k})}\rVert_{1}]\leq\rho_{\pi}$, $\mathbb{E}[\lVert\mu_{i}^{(Q_{k})}-\hat{\mu}_{i}^{(Q_{k})}\rVert_{2}]\leq\rho_{\mu}$, and $\mathbb{E}[d_{\text{Bures}}(\Sigma_{i}^{(Q_{k})},\hat{\Sigma}_{i}^{(Q_{k})})]\leq\rho_{\Sigma}$, $\forall i$ and $\forall k$.
Here the expectations are taken with respect to samples drawn from $Q_{k}$. Under these conditions we have,
Theorem 3.2.
Let $P$ and $Q$ be two labeled GMMs with bounded parameters and satisfying the empirical convergence rates. Then, the following holds,
$\mathbb{E}[\hat{\mathbb{B}}(P_{\tau})-\hat{\mathbb{B}}^{\star}]\leq e^{-C_{\text{PL}\tau}}(\mathbb{B}(P_{0})-\mathbb{B}^{\star})+C_{\lambda,\mathcal{Q}},$
where $C_{\lambda,\mathcal{Q}}$ is a constant that depends on the coordinates $\lambda$ and GMMs in $\mathcal{Q}$.
Here $C_{\lambda,\mathcal{Q}}$ depends on the $(\pi^{(Q_{k})},\mu^{(Q_{k})},\Sigma^{(Q_{k})})$ of each GMM, but not on $\nu^{(Q_{k})}$. Indeed, since we fit a GMM per class, the labels are estimated exactly (i.e., $P_{i,x}$ and $\hat{P}_{i,x}$ belong to the same class).
@@ 4 Experiments
We divide our experiments into a toy example (Section 4.1) to illustrate our methods, and multi-source domain adaptation experiments (Section 4.2). Due to length constraints, we include additional experiments in Sections 5 and 6 in the Appendix. In particular, in Table 1 in the Appendix, we show a running-time comparison demonstrating the scalability of our method.
4.1 Toy Example
In Figure 2, we show the Swiss roll measure $Q_{0}$, alongside four variations obtained via $T_{k,\sharp}Q_{0}$, $k=1,\cdots,4$. This example has been used in several state-of-the-art Wasserstein barycenter works [korotin2021continuous, visentin2025computing]. Originally there are labels associated with the samples of the Swiss-roll measure, corresponding to the position of the points in the underlying manifold. We show a summary of our results in Figure 3.
Figure 2: Location-scatter family generated by a Swiss-roll measure, $Q_{0}$. Each measure $Q_{k}=T_{k,\sharp}Q_{0}$, for $T_{k}(x)=A_{k}x+b_{k}$.
First, we compare unsupervised barycenter solvers in Figures 3 (a) through (h). Quantitatively, empirical methods, notably [cuturi2014fast, Algorithm 2] and our WGF algorithm (cf. Algorithm 1) achieve the lowest Wasserstein distance to the ground truth in Figure 2. We note that, overall, while the neural network solvers are usually more scalable than discrete methods in terms of number of samples, their optimization problem is more complicated and very sensitive to hyper-parameters.
(a) CW2B [korotin2021continuous].
(b) NOT [kolesov2024estimating].
(c) Discrete [cuturi2014fast].
(d) WGF.
@@ 4.1 Toy Example
(g) U-NOT [gazdieva2024robust].
(h) NormFlow [visentin2025computing].
(i) Discrete [montesuma2023multi].
(j) WGF.
@@
Figure 3: Comparison between Wasserstein barycenter solvers. Colored scatter plots indicate labeled barycenters. For each solver, we compute the Wasserstein distance between its solution $\hat{P}$ and the ground-truth $P^{\star}$, shown in the title of each sub figure (best seen on screen). Overall, using label information leads to barycenters that better approximate the ground-truth barycenter.
Second, we experiment with integrating labels in the ground cost as described in Section 3.3. These methods are shown in Figures 3 (i) through (l). In all cases, integrating labels produces barycenters that are closer to the ground-truth. We conclude that using the labels gives a strong inductive bias in the barycenter computation, which helps explain the gain in performance of labeled barycenters in the next section.
4.2 Multi-Source Domain Adaptation
Problem Formulation. One of the main applications of Wasserstein barycenters is multi-source domain adaptation (MSDA) [kouw2019review, sun2015survey, pan2009survey]. In this setting one needs to adapt multiple labeled source measures $Q_{1},\cdots,Q_{K}$ to a single unlabeled target measure $Q_{T}$. The goal is to learn, from samples $\{\{x_{i}^{(Q_{k})},y_{i}^{(Q_{k})}\}_{i=1}^{n_{k}}\}_{k=1}^{K}$ and $\{x_{i}^{(Q_{T})}\}_{i=1}^{n_{T}}$, a classifier $h$ that achieves low risk in the target domain measure,
$\mathcal{R}_{Q_{T}}(h)=\mathbb{E}_{(x,y)\sim Q_{T}}[\mathcal{L}(y,h(x))],$
for a loss function $\mathcal{L}$ (e.g., cross-entropy loss). We isolate the quality of barycenters by doing adaptation at the level of embeddings. This approach allows us to perform domain adaptation in a higher semantic space, where distributions are more meaningful and comparable across domains. Thus, we assume that a meaningful feature extractor $\phi$, called the backbone, has been previously learned. We obtain the feature extractor by fine-tuning a neural network on the labeled source domain data. In addition, similar to [montesuma2021wasserstein, montesuma2024lighter], we align the barycenter with the target through OT [courty2016optimal]. More details are available in Section 6 in the Appendix.
Benchmark Backbone # Samples # Domains # Dim. # Classes
Office31 ResNet50 [he2016deep] 3287 3 2048 31
BCI-CIV-2a CBraMod [wang2024cbramod] 5184 10 200 4
TEP CNN [montesuma2024benchmarking] 17289 6 128 29
Office Home ResNet101 [he2016deep] 15500 4 2048 65
ISRUC CBraMod [wang2024cbramod] 89240 100 512 5
Experimental Setting. We run our experiments on five benchmarks: Office 31 [saenko2010adapting], Office Home [venkateswara2017deep], BCI-CIV-2a [brunner2008bci], ISRUC [khalighi2013automatic], and TEP [montesuma2024benchmarking]. The first two, second two, and last benchmarks correspond to computer vision, neuroscience, and chemical engineering benchmarks, respectively. We show in Table 1 an overview of our experimental setting.
Compared Methods. Overall, we compare seven Wasserstein barycenter strategies with ours. These methods are: discrete barycenters [cuturi2014fast, Algorithm 2], normalizing flows [visentin2025computing], continuous 2-Wasserstein barycenters [korotin2021continuous], neural OT barycenters [kolesov2024estimating], and unbalanced neural OT barycenters [gazdieva2024robust]. We also include labeled barycenter strategies [montesuma2023multi, montesuma2024multi]. A brief review of these methods is available in Section 2 in the Appendix. For completeness, we include four other state-of-the-art methods in domain adaptation over embedding vectors. Those methods are: WJDOT [turrisi2022multi], DaDiL-R and E [montesuma2023multi], and GMM-DaDiL [montesuma2024lighter]. For each benchmark, we use one domain as the target domain (e.g., Amazon vs. $\{\text{dSLR, Webcam}\}$ in the Office 31 benchamrk), and we measure the classification accuracy, i.e., the percentage of correct predictions.
Benchmark $\mathcal{X}\times\mathcal{Y}$ Office31 OfficeHome BCI-CIV-2a ISRUC TEP
Backbone - ResNet-50 ResNet-101 CBraMod CBraMod CNN
Source-Only - 86.40 75.95 50.30 76.63 78.48
WJDOT [turrisi2022multi] - 86.80 76.59 N/A 76.95 86.13
DaDiL-R [montesuma2023multi] - 89.91 77.86 53.41 74.68 86.14
DaDiL-E [montesuma2023multi] - 89.79 78.14 N/A 75.89 85.87
GMM-DaDiL [montesuma2024lighter] - 90.63 78.81 57.10 75.47 86.85
Discrete [cuturi2014fast] 81.94 70.42 57.38 70.67 83.81
NormFlow [visentin2025computing] 85.91 76.73 55.45 78.04 82.89
CW2B [korotin2021continuous] 86.37 76.44 57.25 75.84 85.83
NOT [kolesov2024estimating] 86.22 75.36 57.29 76.39 84.93
U-NOT [gazdieva2024robust] 86.97 76.85 57.17 77.06 85.43
Discrete [montesuma2023multi] 87.93 77.09 57.43 78.20 86.09
GMM [montesuma2024lighter] 88.54 77.87 57.04 74.58 84.67
WGF (ours) 88.14 77.83 57.50 79.78 86.21
WGF-GMM (ours) 89.31 78.71 57.35 78.72 86.77
Table 2: Average classification accuracy on target domains. In total, we compare eight methods against our WGF strategy. For barycenter solvers, we indicate by ✓methods that compute barycenters on $\Omega=\mathcal{X}\times\mathcal{Y}$. Bold numbers represent the best barycenter method, and underlined numbers represent second best.
Main results. We present our main results in Table 2, which reports the average performance per domain on each benchmark. We provide fine-grained results in Section 6 in the Appendix. Overall, labeled barycenter methods (the four last rows in Table 2) have a clear advantage with respect to the unsupervised methods. We argue that label information is pivotal in domain adaptation success, which is consistent with previous research [courty2016optimal, montesuma2021wasserstein, montesuma2023multi, montesuma2024lighter]. Among the labeled barycenter methods, our gradient flow framework achieves either the best performance, or remains competitive, surpassing previous methods in MSDA, such as WJDOT and DaDiL in the ISRUC benchmark.
Figure 4: t-SNE [maaten2008visualizing] visualization of the barycenter of $\text{Art},\text{Product},\text{Real-World}]$ source domains in the Office home benchmark. Colors represent different classes from $1$ to $65$. Overall, the classes tend to be more separated when using the repulsion functional $\mathbb{U}$.
Visualization. We now visualize the effect of adding the repulsion interaction energy functional on the Office-Home benchmark. Adding $\mathbb{U}$ leads to classes that are well separated. For computer vision experiments we use ResNets [he2016deep] as the backbone, which produce $2048$-dimensional embeddings. For that reason, we use the cosine distance $d(x,x^{\prime})=1-\text{cossim}(x,x^{\prime})$ in equation 21 and we fix $\text{margin}=1$. This choice encourages class clusters to live in orthogonal sub-spaces of the latent space. We show a comparison of the obtained barycenters in Figure 4. Overall, the classes tend to have less overlap when the interaction energy is added to the barycenter functional.
Ablation. We ablate the effect of $\mathbb{V}$ and $\mathbb{U}$ in the Office Home benchmark. For this benchmark, we use $\mathbb{V}(P)=\mathbb{W}_{2}(\hat{P},\hat{Q}_{T})^{2}$. The intent here is to isolate the contribution of each component to the state-of-the-art results in Table 2. Our results are summarized in Table 3, where we further compare our results with other related barycenter methods. In the empirical setting, the use of Algorithm 1 alone already presents a gain in performance compared to its discrete counterpart. Overall, the combination of $\mathbb{B}+\mathbb{V}+\mathbb{U}$ consistently gives the best performance, showing the importance of adding underlying structure to the Wasserstein barycenter in domain adaptation.
Method $\mathcal{X}\times\mathcal{Y}$ $\mathbb{B}$ $\mathbb{B}+\mathbb{V}$ $\mathbb{B}+\mathbb{U}$ $\mathbb{B}+\mathbb{V}+\mathbb{U}$
Discrete [cuturi2014fast] 70.42 N/A N/A N/A
Discrete [montesuma2023multi] 77.09 N/A N/A N/A
WGF (ours) 77.48 77.07 77.08 77.83
GMM [montesuma2024lighter] 77.87 N/A N/A N/A
WGF-GMM (ours) 77.28 78.28 78.12 78.71
Table 3: Ablation of average target domain classification accuracy on the Office Home benchmark.

### v3

@@ 3.4 Convergence
One of the difficulties in analyzing the gradient flow of the functional in equation 12 comes from the fact that $P\mapsto\mathbb{W}_{2}(P,Q)^{2}$ is not geodesically convex in $\mathcal{P}_{2}(\Omega)$ [santambrogio2017euclidean, Section 4.4]. As a result, we are minimizing a non-convex functional. Explicit constants and proofs are available in Appendix D. Our analysis relies on a measure-theoretic version of the Polyak Łojasiewicz (PL) inequality [polyak1964gradient, karimi2016linear],
$\lVert\nabla_{W}\mathbb{B}(P)\rVert_{L_{2}(P)}^{2}\geq C_{\text{PL}}(\mathbb{B}(P)-\mathbb{B}^{\star}).$ (23)
When this inequality holds, we have the following convergence result.
@@
Let $\Omega=\mathcal{B}(0,R)$ be the closed ball in $\mathbb{R}^{d}$ with radius $R>0$. Let $P^{\star}\in\mathcal{P}_{2}(\Omega)$ be the barycenter of $\mathcal{Q}=\{Q_{k}\}_{k=1}^{K},Q_{k}\in\mathcal{P}_{2}(\Omega)$ $k=1,\cdots,K$, with barycentric coordinates $\lambda=(\lambda_{1},\cdots,\lambda_{K})\in\Delta_{K}$. Approximate each $Q_{k}$ with $m$ i.i.d. samples $\{z_{i}^{(Q_{k})}\}_{i=1}^{m}$, $z_{i}^{(Q_{k})}\sim Q_{k}$ and $\{z_{\tau,i}^{(P)}\}_{i=1}^{n}$ be the barycenter support. Under inequality 42 the following holds,
$\begin{aligned} \mathbb{E}[\hat{\mathbb{B}}(\hat{P}_{\tau})-\hat{\mathbb{B}}^{\star}]\leq e^{-C_{\text{PL}}\tau}(\mathbb{B}(\hat{P}_{0})-\mathbb{B}^{\star})+C_{R}\sqrt{\dfrac{C_{d,n}}{m}}\end{aligned},$ (24)
where the expectation is taken with respect to samples from $Q_{k}$. The constants $C_{R}$ and $C_{d,n}$ depend on the radius $R$, number of barycenter samples $n$, and dimensions $d$. $m$ denotes the number of samples drawn from each $Q_{k}.$
The first term in the r.h.s. of equation 24 decays exponentially with the iterations $\tau$, and is a direct consequence of the PL inequality in equation 23. The second term captures the empirical approximation error of the involved measures, that is, $Q_{k}\in\mathcal{Q}$ and $P_{\tau}$. In practice, $\tau\rightarrow+\infty$ leads to an error governed by the empirical approximations.
On the PL Inequality. Equation 23 is a standard tool used in proving the convergence of non-convex optimization algorithms, however, it is an open question whether it holds in general settings. For instance, chewi2020gradient shows that this inequality holds for Guassian measures under specific spectral conditions. In the Appendix, we show that the same result holds for location-scatter families of measures (Appendix D.2), based on the isometry between these sets, and the Bures-Wasserstein manifold established by [alvarez2016fixed, Theorem 2.3]. On top of this theoretical justification, we provide an experiment validating the bound on the Swiss roll measure (Appendix D.1).
(a) $Q_{0}$.
(b) $Q_{1}$.
(c) $Q_{2}$.
(d) $Q_{3}$.
(e) $Q_{4}$.
(f) $P^{\star}$.
Figure 2: Swiss roll measures in a location-scatter family, where a source measure is transformed under an affine function $T_{k}(x)=A_{k}x+b_{k}$. Colors reflect the relative position of points in the manifold.
@@ 4 Experiments
In the following, all experiments were conducted on a virtual machine with 8 CPUs (AMD EPYC 7413, 48GB of RAM) and an NVIDIA L4 GPU (24GB of VRAM). Appendix F includes additional details, fine-grained results, complexity analysis and ablations.
4.1 Swiss Roll Measures
In Figure 2, we show the Swiss roll measure [korotin2021continuous] $Q_{0}$, alongside four variations obtained via $T_{k,\sharp}Q_{0}$, $k=1,\cdots,4$, $T_{k}(x)=A_{k}x+b_{k}$, for which the ground-truth barycenter is known in closed-form [alvarez2016fixed], that is, $P^{\star}=(\sum_{k}\lambda_{k}T_{k})_{\sharp}Q_{0}$. These measures are shown in Figure 2.
(a) Discrete.
(b) CW2B.
(c) U-NOT.
(d) NormFlow.
(e) WGF.
@@ 4.1 Swiss Roll Measures
Figure 3: Comparison of unsupervised barycenter methods. Solvers (b, c, d) are neural, and (e, f) are ours.
We compare 5 barycenter solvers, including discrete solvers [cuturi2014fast, montesuma2023multi], and neural solvers: CW2B [korotin2021continuous], U-NOT gazdieva2024robust, and NormFlow visentin2025computing. First, we benchmark unsupervised solvers in Figures 3 (a) through (f). Quantitatively, empirical methods, notably [cuturi2014fast, Algorithm 2] and our Wasserstein Gradient Flow (WGF) algorithm (cf. Algorithm 1) achieve the lowest Wasserstein distance to the ground truth in Figure 2. Overall, while the neural network solvers are usually more scalable than discrete methods in terms of number of samples, their optimization is more complicated and very sensitive to hyper-parameters.
Second, we experiment with integrating labels in the ground cost as described in Section 3.2. These methods are shown in Figures 4 (a), (b) and (c). In all cases, integrating labels produces barycenters that are closer to the ground-truth. We conclude that using the labels gives a strong inductive bias in the barycenter computation, which explains the gain in performance of labeled barycenters in the next section.
(a) Discrete.
(b) WGF.
@@
Figure 4: Comparison of supervised barycenter methods.
Next, we compare the running time of our WGF solver with the discrete solver of cuturi2014fast, for an increasing barycenter support size $n\in\{2^{10},\cdots,2^{16}\}$ and batch size $m\in\{2^{8},2^{9},2^{10}\}$ for $200$ iterations. The discrete solver is run for $n\in\{2^{10},\cdots,2^{14}\}$. For this algorithm, larger support sizes results in memory overflow. Our WGF is able to compute larger barycenters via mini-batching.
We summarize our results in Figure 5. First, in Figure 5 (a), our WGF leads to speedups ranging from $2\times$ $(n=2^{12},m=2^{10})$ to $50\times$ $(n=2^{14},m=2^{8})$. Similar gains were obtained using Exact OT on CPU (Appendix F.3). Second, in Figure 5 (b), we isolate the gain of using GPU acceleration and our vectorization strategy (Section 3.1), showing consistent gains ranging from $2.1\times$ to $8.3\times$. Overall, our WGF strategy provides compounding speedups from three sources: (i) minibatching, (ii) entropic regularization, and (iii) GPU parallelism.
(a) Scaling with $m$ ($\epsilon>0$)
(b) Scaling with $\epsilon$ ($m=256$)
Figure 5: Running time analysis for our WGF algorithm. In (a), we compare the running time of our technique in comparison with the discrete solver of cuturi2014fast. In (b), we isolate the effect of GPU acceleration, showing consitent gains with increasing support size.
4.2 Multi-Source Domain Adaptation
One of the main applications of Wasserstein barycenters is multi-source domain adaptation (MSDA) [kouw2019review, sun2015survey, pan2009survey]. In this setting one needs to adapt multiple labeled source measures $Q_{1},\cdots,Q_{K}$ to a single unlabeled target measure $Q_{T}$. The goal is to learn, from samples $\{\{x_{i}^{(Q_{k})},y_{i}^{(Q_{k})}\}_{i=1}^{n_{k}}\}_{k=1}^{K}$ and $\{x_{i}^{(Q_{T})}\}_{i=1}^{n_{T}}$, a classifier $h$ that achieves low risk or error in the target domain measure,
$\mathcal{R}_{Q_{T}}(h)=\mathbb{E}_{(x,y)\sim Q_{T}}[\mathcal{L}(y,h(x))],$
for a loss function $\mathcal{L}$ (e.g., cross-entropy loss). We isolate the quality of barycenters by doing adaptation at the level of embeddings. This approach allows us to perform domain adaptation in a higher semantic space, where distributions are more meaningful and comparable across domains. Thus, we assume that a meaningful feature extractor $\phi$, called the backbone, has been previously learned. We obtain the feature extractor by fine-tuning a neural network on the labeled source domain data (Appendix E.2).
Experimental Setup. The unsupervised WGF is fit with unlabeled measures, i.e., $Q_{k}=n_{k}^{-1}\sum_{i}\delta_{x_{i}^{(Q_{k})}}$, where $x_{i}^{(Q_{k})}$ are embedding vectors. The supervised WGF is fit with labeled measures, that is, $Q_{k}=n_{k}^{-1}\sum_{i=1}^{n_{k}}\delta_{(x_{i}^{(Q_{k})},y_{i}^{(Q_{k})})}$. In both cases, $\hat{P}=n^{-1}\sum_{i=1}^{n}\delta_{(x_{i}^{(P)},y_{i}^{(P)})}$ which allows us to apply the regularizing functionals defined in section 3.3.
Wasserstein Barycenter Transport. montesuma2021wasserstein proposed a technique based on Wasserstein barycenters for domain adaptation. The idea is to map the synthesized barycenter support to the target domain through the barycentric mapping [courty2016optimal]. Then, a classifier can be learned on the target domain based on the transported data. We adapt this idea to unsupervised barycenter algorithms by using the computed barycenter as a pivot domain. See appendix E.4 for more details.
Benchmarks. We run our experiments on five benchmarks: Office 31 [saenko2010adapting], Office Home [venkateswara2017deep], BCI-CIV-2a [brunner2008bci], ISRUC [khalighi2013automatic], and TEP [montesuma2024benchmarking]. The first two, second two, and last benchmarks correspond to computer vision, neuroscience, and chemical engineering benchmarks, respectively. We show in Table 1 an overview of our experimental setting.
Benchmark Backbone # Samples # Domains # Dim. # Classes
Office31 ResNet50 3287 3 2048 31
BCI-CIV-2a CBraMod 5184 10 200 4
TEP CNN 17289 6 128 29
Office Home ResNet101 15500 4 2048 65
ISRUC CBraMod 89240 100 512 5
Compared Methods. Overall, we compare seven Wasserstein barycenter strategies with ours. Besides those used in the previous section, we include the NOT approach of kolesov2024estimating, and the Gaussian Mixture Model (GMM) barycenter of montesuma2024lighter. For completeness, we include four other state-of-the-art methods in domain adaptation over embedding vectors. Those methods are: WJDOT turrisi2022multi, DaDiL-R, and E montesuma2023multi, and GMM-DaDiL montesuma2024lighter. For each benchmark, we use one domain as the target domain (e.g., Amazon vs. $\{\text{dSLR, Webcam}\}$ in the Office 31 benchmark), and we measure the classification accuracy, i.e., the percentage of correct predictions.
Main Results. We present our main results in Table 2, which reports the average performance per domain on each benchmark. We provide fine-grained results in Appendix F. In general, labeled barycenter methods (the 3 last rows in Table 2) have a clear advantage over unsupervised methods. We argue that label information is essential in domain adaptation success, which is consistent with previous research [courty2016optimal, montesuma2021wasserstein, montesuma2023multi, montesuma2024lighter]. Among barycenter methods, our WGF method achieves the best performance in all benchmarks, outperforming previous methods in MSDA, such as WJDOT and DaDiL in the ISRUC, BCI-CIV, and TEP benchmarks.
Benchmark $\mathcal{X}\times\mathcal{Y}$ Office31 OfficeHome BCI-CIV-2a ISRUC TEP Avg. Rank
Backbone - ResNet-50 ResNet-101 CBraMod CBraMod CNN -
Source-Only - 86.40 75.95 50.30 76.63 78.48 -
WJDOT - 86.80 76.59 N/A 76.95 86.13 7.8
DaDiL-R - 89.91 77.86 53.41 74.68 86.14 6.4
DaDiL-E - 89.79 78.14 N/A 75.89 85.87 5.0
GMM-DaDiL - 90.63 78.81 57.10 75.47 86.85 4.4
Discrete 87.82 77.17 57.38 70.67 83.81 8.8
NormFlow 85.91 76.73 55.45 78.04 82.89 10.0
CW2B 86.37 76.44 57.25 75.84 85.83 9.0
NOT 86.22 75.36 57.29 76.39 84.93 9.4
U-NOT 86.97 76.85 57.17 77.06 85.43 7.8
@@ 4.2 Multi-Source Domain Adaptation
Table 2: Average classification accuracy on target domains. The top group (WJDOT through GMM-DaDiL) include general domain adaptation methods. The middle and bottom groups show barycenter solvers without and with access to source labels ($\mathcal{X}\times\mathcal{Y}=$✗, $\mathcal{X}\times\mathcal{Y}=$✓), respectively. Bold indicates best among barycenter methods.
Unsupervised vs. Supervised Barycenters. An intriguing finding of Table 2 is that, on some benchmarks, unsupervised WGF achieves decent performance, even outperforming supervised baselines (e.g., WJDOT of turrisi2022multi). We probe the reason why these methods have unreasonably good performance, by performing a t-SNE [maaten2008visualizing] visualization of the barycenter against the source domain samples, shown in Figure 6. We see that even without labels, the barycenter samples cluster around the classes. This effect stems from the geometry of the embeddings, which were optimized (through supervision on source domains) to separate classes. As a result, unsupervised OT already captures, to some extent, the cluster structure of the barycentric measure. Figures 6 (b) and (c) investigate the remaining performance gap, as we transport the measures towards the target. Figure 6 (b) shows the transported sources to the barycentric domain (left) and the synthesized barycenter (right), and Figure 6 shows the final transported samples in the target domain for the unsupervised strategy (left) and supervised one (right). We conclude that using a synthesized supervised barycenter is better than using an unsupervised barycenter as pivot domain (Appendix E.4). Therefore, the inductive bias of using labels in the ground cost is essential for domain adaptation. This further explains the performance of DaDiL methods, which directly synthesize the target domain through an approximate barycenter.
(a) Unsupervised vs. Supervised barycenters.
(b) Transported sources vs. supervised barycenter.
(c) Transported samples in the target domain.
Figure 6: Comparison of unsupervised vs. supervised barycenter transports on the TEP benchmark.